Section 13

Combinatorics

In case of several authors, Invited Speakers are marked with a ^{∗}.

Béla Bollobás: Hereditary Properties of Graphs: Asymptotic Enumeration, Global Structure, and Colouring . . . . . . . . . . . . . . . . . . . . . III 333

András Frank: Applications of Relaxed Submodularity . . . . . . . . . . . . III 343

Alain Lascoux: Ordonner le Groupe Symétrique: Pourquoi Utiliser l’Algèbre de Iwahori-Hecke ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 355

Jiří Matoušek: Mathematical Snapshots from the Computational Geometry Landscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 365

Harald Niederreiter: Nets, (t, s)-Sequences, and Algebraic Curves over Finite Fields with Many Rational Points . . . . . . . . . . . . . . . . . . . . . . III 377

N. J. A. Sloane: The Sphere Packing Problem . . . . . . . . . . . . . . . . . . . . . III 387
Joseph A. Thas: Finite Geometries, Varieties and Codes . . . . . . . . . . . III 397

Andrei Zelevinsky: Multisegment Duality, Canonical Bases and Total Positivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 409

Documenta Mathematica · Extra Volume ICM 1998 · III ·

332

Documenta Mathematica · Extra Volume ICM 1998 · III ·